# Canonical Autoregressive Generation

### v1

@@ 3 An Efficient Approach to Canonical Autoregressive Generation
In this section, we start by showing that, for an output token sequence generated by an LLM to be canonical, the partial token sequences generated at each step of the autoregressive generation must also be canonical. Building upon this result, we then introduce canonical sampling, a next-token sampling method that enables an LLM to generate only canonical token sequences, as well as an efficient implementation of canonical sampling using the Gumbel-Max trick [35]. Finally, we conclude by showing that, in comparison with standard sampling, the distribution of token sequences generated using canonical sampling is provably closer to the true distribution of token sequences used during training.
@@ 3.1 Subsequences of Canonical Token Sequences Must Also Be Canonical
Our starting point is the following theorem, which tells us that, if an LLM generates a non-canonical (partial) output sequence at any step of the autoregressive generation process, the output sequence is bound to remain non-canonical in subsequent steps: [^3] All proofs can be found in Appendix A.
@@
Let ${\mathcal{T}}=(\Sigma,V,\texttt{enc},\texttt{dec})$ be a BPE-based tokenizer [^4] Our results generalize to Unigram- and Wordpiece-based tokenizers, which we discuss further in Appendix B. and $\mathbf{s}\in V^{*}$ a non-canonical token sequence according to ${\mathcal{T}}$. Then, for any token $t\in V$, $\mathbf{s}\shortmid t$ is also non-canonical.
Proof sketch.
We prove the theorem by contradiction. We first assume that there exists a non-canonical sequence $\mathbf{s}$ and a token $t$ such that $\mathbf{s}\shortmid t$ is canonical. This implies that, applying in order the merge rules of the tokenizer’s merge rule sequence $\mathcal{R}$, to the single-character token sequence corresponding to the string $\texttt{dec}(\mathbf{s}\shortmid t)$, results in the token sequence $\mathbf{s}\shortmid t$ itself. We then show that the applied merges can be split into two separate groups that would independently result in $\mathbf{s}$ and $t$ starting from the corresponding strings $\texttt{dec}(\mathbf{s})$ and $\texttt{dec}(t)$, respectively. Based on that, we argue that, since $\mathbf{s}$ is non-canonical, there must exist a different, canonical tokenization $\mathbf{s}^{\prime}\neq\mathbf{s}$ of the string $\texttt{dec}(\mathbf{s})$, created by applying a different sequence of merge rules than the one that resulted in $\mathbf{s}$. The rest of the proof is based on the observation that, since the merges that create the last token $t$ while transforming $\texttt{dec}(\mathbf{s}\shortmid t)$ into $\mathbf{s}\shortmid t$ are independent of the merges that transform $\texttt{dec}(\mathbf{s})$ into $\mathbf{s}$, the string $\texttt{dec}(\mathbf{s})$ should have been tokenized following the merge rules that result in $\mathbf{s}^{\prime}$, which contradicts the fact that $\mathbf{s}\shortmid t$ is canonical. ∎
The above theorem readily implies that an output token sequence generated by an LLM is canonical if and only if the partial token sequences generated at each step of the autoregressive generation process are canonical. This theorem also provides a plausible explanation to the empirical observation that the likelihood that an LLM generates non-canonical output sequences increases with the length of the sequence [6]. This is because, since sampling a “non-canonical token” once during the autoregressive process is sufficient to render the output token sequence non-canonical, it is natural that the chances of this to happen increase with the number of sampled tokens.
3.2 Canonical Sampling
At each step of the autoregressive generation process, canonical sampling sets the probability of a subset of tokens to zero—those that, when appended to a partial output sequence, would result in a non-canonical token sequence—and redistributes their probability mass to the remaining tokens proportionally to their original probability mass.
Formally, let $d_{\mathbf{s}}$ denote the next-token distribution generated by the LLM given a partial output token sequence $\mathbf{s}$ and let $d_{\mathbf{s}}(t)$ denote the probability of sampling a token $t$ from this distribution. Given the partial output token sequence $\mathbf{s}$, an LLM using canonical sampling draws the next token in the autoregressive generation process from a canonicalized next-token distribution
@@ 3.2 Canonical Sampling
where $Z=\sum_{t\in V:\ {\mathbf{s}}\shortmid t\text{ is canonical}}d_{\mathbf{s}}(t)$ is a normalization constant that ensures that $\tilde{d}_{\mathbf{s}}$ is a valid probability distribution. In that context, note that redistributing the probability mass of tokens that would lead to non-canonical token sequences proportionally to the original probabilities $d_{\mathbf{s}}(t)$ is a natural choice we make, inspired by other popular sampling methods used by LLMs, such as top-$k$ and top-$p$ sampling [36].
To compute the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$ and sample from it directly, one would need to evaluate whether the token sequence ${\mathbf{s}}\shortmid t$ is canonical for every token $t\in V$, which can be computationally expensive, especially since these evaluations are required at every step of the autoregressive generation process. Fortunately, we can efficiently sample from the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$ using Algorithm 1.
Algorithm 1 starts by sampling a value $u_{t}\sim\text{Gumbel}(0,1)$ from a Gumbel distribution for each token $t\in V$. Then, it ranks the tokens in decreasing order with respect to the perturbed log-probability $\log(d_{\mathbf{s}}(t))+u_{t}$. Finally, it returns the token $t$ with the largest value of $\log(d_{\mathbf{s}}(t))+u_{t}$ such that ${\mathbf{s}}\shortmid t$ is canonical. Algorithm 1 relies on a property of the Gumbel-Max trick [37, 35], which states that the $\mathop{\mathrm{argmax}}$ operation over a constrained subset of categorical outcomes is equivalent to sampling from a categorical distribution with zero probability for all outcomes outside the subset, and with the probabilities of the outcomes in the subset scaled proportionally to their original probabilities, as shown in Eq. 2 in Maddison et al. [37]. Hence, it readily holds that Algorithm 1 returns a valid sample from the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$ defined in Eq. 1, i.e.,
@@
Algorithm 1 Canonical Sampling with Gumbel-Max
@@
Further, it is worth highlighting that, in contrast to computing the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$, which requires evaluating the canonicity of $|V|$ token sequences, Algorithm 1 requires only a few evaluations of canonicity. This is because, in practice, LLMs tend to generate mostly canonical token sequences [6], hence, the probabilities $d_{\mathbf{s}}(t)$ generated by an LLM for tokens $t$ that lead to non-canonical sequences ${\mathbf{s}}\shortmid t$ are usually small. More specifically, let $d_{\mathbf{s}}(\text{canonical})$ be the probability mass on the subset of tokens that lead to canonical sequences, i.e., $d_{\mathbf{s}}(\text{canonical})=\sum_{t\in V:\ {\mathbf{s}}\shortmid t\text{ is canonical}}d_{\mathbf{s}}(t)$, then Algorithm 1 requires, in expectation, fewer than $1/d_{\mathbf{s}}(\text{canonical})$ evaluations of canonicity before successfully sampling the next token. That is because, unlike (independent) rejection sampling from $d_{\mathbf{s}}$, which would require in expectation exactly $1/d_{\mathbf{s}}(\text{canonical})$ evaluations of canonicity until a token that leads to a canonical sequence is successfully sampled, our approach never checks the same token twice, which results in an increase in the success probability of sampling a token that leads to a canonical sequence after each failed attempt. [^5] The number of evaluations of canonicity in rejection sampling is distributed according to a geometric distribution with success probability $d_{\mathbf{s}}(\text{canonical})$ resulting in $1/d_{\mathbf{s}}(\text{canonical})$ evaluations in expectation until a successful sample.
3.3 Canonical Sampling Brings us Closer to the True Distribution of Token Sequences
@@
Let $p$ denote the true distribution over token sequences ${\mathbf{s}}\in V^{*}$ in the (token) language used during training, for which note that it holds that $p(\mathbf{s})=0$ for all sequences $\mathbf{s}$ that are non-canonical. Moreover, let $d$ denote the distribution over token sequences that the LLM generates using standard sampling, and $\tilde{d}$ the distribution over token sequences that the LLM generates using canonical sampling, that is, sampling from the canonicalized next token distribution $\tilde{d}_{\mathbf{s}}$ given by Eq. 1 at each step of the autoregressive generation process. Then, the following theorem shows that $p$ is likely to be closer to $\tilde{d}$ than $d$ in terms of KL-divergence:
@@
Let $d$ be absolutely continuous [^6] Absolute continuity is required for the KL-divergence to be well defined, i.e., we require that $d(\mathbf{s})=0$ implies that $p(\mathbf{s})=0$ for all $\mathbf{s}\in V^{*}$. with respect to $p$. Moreover, assume that there exist $\mathbf{s}\in V^{*}$ and $t_{1},t_{2}\in V$ such that $\mathbf{s}\shortmid t_{1}$ is non-canonical with $d(\mathbf{s}\shortmid t_{1})>0$ and $\mathbf{s}\shortmid t_{2}$ is canonical with $p(\mathbf{s}\shortmid t_{2})>0$ and $d(\mathbf{s}\shortmid t_{2})>0$. Then, it holds that
@@
In words, the two assumptions for the theorem to hold are that (i) there must exist non-canonical token sequences with positive probability of being generated under $d$ so that their probability mass can be redistributed, and (ii) there must exist canonical token sequences with positive probability under $d$ and $p$ so that the redistribution of probability mass in $\tilde{d}$ is beneficial. To understand the intuition behind Theorem 2, note that, by using canonical sampling (i.e., sampling from $\tilde{d}$ instead of $d$), the probability that an LLM generates non-canonical token sequences becomes zero, and the probability that it generates any other (canonical) token sequence increases under $\tilde{d}$. Further, since only canonical token sequences have positive probability under the true distribution $p$, this redistribution of probability mass from non-canonical token sequences to canonical ones can only bring the distribution $\tilde{d}$ closer to the true distribution $p$ compared to $d$.
@@
On the flip side, it is important to clarify that a similar property does not necessarily hold for the respective distributions over strings. That is, using canonical sampling, the distribution of output strings, resulting from decoding the output token sequences, is not guaranteed to be closer (in terms of KL-divergence) to the true distribution of output strings used during training. Formally, let $p_{\texttt{dec}}=P_{\mathbf{s}\sim p(\mathbf{s})}[\texttt{dec}(\mathbf{s})]$ be the true distribution over strings, $d_{\texttt{dec}}=P_{\mathbf{s}\sim d(\mathbf{s})}[\texttt{dec}(\mathbf{s})]$ be the distribution of strings induced by the distribution of output token sequences $d$, and $\tilde{d}_{\texttt{dec}}=P_{\mathbf{s}\sim\tilde{d}(\mathbf{s})}[\texttt{dec}(\mathbf{s})]$ be the distribution of strings induced by the distribution of output token sequences $\tilde{d}$. Then, we can show that
$\textrm{KL}(p_{\texttt{dec}},\tilde{d}_{\texttt{dec}})=\textrm{KL}(p,\tilde{d})$
@@
because there is a one-to-one mapping determined by the encoder enc from any string to a canonical token sequence, and only canonical token sequences have positive probability under $p$ and $\tilde{d}$. In contrast, one cannot claim the same for $\textrm{KL}(p_{\texttt{dec}},d_{\texttt{dec}})$ and $\textrm{KL}(p,d)$, as the same string can have multiple tokenizations that have positive probability under $d$. Thus, Theorem 2 does not imply that $\textrm{KL}(p_{\texttt{dec}},\tilde{d}_{\texttt{dec}})<\textrm{KL}(p_{\texttt{dec}},d_{\texttt{dec}})$.

### v2

@@ 4 Avoiding Tokenization Multiplicity through Canonical Generation
In what follows, we first show that, for an output token sequence generated by an LLM to be canonical, the partial token sequences generated at each step of generation must also be canonical. Building upon this result, we then introduce an efficient and easy-to-implement sampling algorithm for canonical generation based on the Gumbel-Max trick (Huijben et al., 2022). Finally, we conclude by analyzing both theoretically and empirically the quality of the outputs generated by canonical generation against those generated by standard generation, as well as the runtime of our sampling algorithm.
@@ 4.1 Subsequences of Canonical Token Sequences Must Also Be Canonical
In this section, we establish our main theoretical result, which shows that the most commonly used tokenizers—BPE, Unigram, and Wordpiece—are non-recovering tokenizers.
Definition 1.
A tokenizer ${\mathcal{T}}=(\Sigma,V,\texttt{enc},\texttt{dec})$ is called non-recovering if it holds that, for any non-canonical token sequence $\mathbf{s}\in V^{+}$ according to $\mathcal{T}$} and any token $t\in V$, $\mathbf{s}\shortmid t$ is also non-canonical.
More formally, we have the following theorem: [^8] The proof for Theorem 2 can be found in Appendix B. In Appendix C, we also prove that, under mild conditions, the theorem holds when using pretokenization.
@@
BPE-, Unigram- and Wordpiece-based tokenizers are non-recovering.
Figure 4: Tokenization multiplicity across languages. The plots show the number of inputs prompts for which we observe at least two outputs given by gpt5m with the same string but different tokenization lengths. Each panel corresponds to one of the three tasks we consider in our experiments and pairs of letters on the x-axis correspond to different languages. Refer to Table 2 in Appendix F for details regarding the languages we use and to Appendix G.1 for qualitatively similar results using other models.
This theorem immediately implies that an output token sequence is canonical if and only if the partial token sequences generated at each step of the generation process are canonical. Moreover, this theorem also provides a plausible explanation for the empirical observation that the likelihood that an LLM generates non-canonical output sequences increases with the length of the sequence (Geh et al., 2024). This is because, since sampling a “non-canonical token” once during the generation process is sufficient to render the output token sequence non-canonical, it is natural that the chances of this to happen increase with the number of sampled tokens.
In this context, we find it rather surprising that the above theorem holds for BPE, Unigram and Wordpiece since these (deterministic) tokenizers use fundamentally different tokenization techniques: BPE uses a rule based approach, Unigram maximizes the probability of the tokenization, and Wordpiece greedily encodes the text to minimize the number of tokens.
4.2 An Efficient Sampling Algorithm for Canonical Generation
Building upon Theorem 2, we now introduce canonical generation, along with an efficient sampling algorithm to implement it. The core principle of canonical generation is to ensure that the sampled tokens at all steps of the generation are such that the respective (partial) output remains canonical. To this end, at each step, it sets the probability of a subset of tokens to zero—those that, when appended to a partial output sequence, would result in a non-canonical token sequence—and redistributes their probability mass to the remaining tokens proportionally to their original probability mass.
Formally, let $d_{\mathbf{s}}$ denote the next-token distribution generated by the LLM given a partial output token sequence $\mathbf{s}$ and let $d_{\mathbf{s}}(t)$ denote the probability of sampling a token $t$ from this distribution. Given the partial output token sequence $\mathbf{s}$, an LLM using canonical generation draws the next token in the generation process from a canonicalized next-token distribution
@@ 4.2 An Efficient Sampling Algorithm for Canonical Generation
where $Z=\sum_{t\in V:\ {\mathbf{s}}\shortmid t\text{ is canonical}}d_{\mathbf{s}}(t)$ is a normalization constant that ensures that $\tilde{d}_{\mathbf{s}}$ is a valid probability distribution. In that context, note that redistributing the probability mass of tokens that would lead to non-canonical token sequences proportionally to the original probabilities $d_{\mathbf{s}}(t)$ is a natural choice we make, inspired by other popular strategies for (stochastic) generation, such as top-$k$ and top-$p$ sampling (Holtzman et al., 2020), and constrained generation (Beurer-Kellner et al., 2024).
Next, we introduce an efficient and easy-to-implement algorithm to sample from the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$, which avoids explicitly computing the entire distribution $\tilde{d}_{\mathbf{s}}$. The algorithm starts by sampling a value $u_{t}\sim\text{Gumbel}(0,1)$ from a Gumbel distribution for each token $t\in V$. Then, it ranks the tokens in decreasing order with respect to the perturbed log-probability $\log(d_{\mathbf{s}}(t))+u_{t}$. Finally, it returns the token $t$ with the largest value of $\log(d_{\mathbf{s}}(t))+u_{t}$ such that ${\mathbf{s}}\shortmid t$ is canonical. The overall procedure, summarized in Algorithm 1, relies on a property of the Gumbel-Max trick (Maddison et al., 2014; Huijben et al., 2022), which states that the $\mathop{\mathrm{argmax}}$ operation over a constrained subset of categorical outcomes is equivalent to sampling from a categorical distribution with zero probability for all outcomes outside the subset, and with the probabilities of the outcomes in the subset scaled proportionally to their original probabilities, as shown in Eq. 2 in Maddison et al. (2014). Hence, it readily holds that Algorithm 1 returns a valid sample from the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$ defined in Eq. 1, i.e.,
@@
Algorithm 1 Canonical Generation via Gumbel-Max Sampling
@@
Further, it is worth highlighting that, in contrast to computing the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$, which requires evaluating the canonicity of $|V|$ token sequences, Algorithm 1 requires only a few evaluations of canonicity. This is because, in practice, LLMs tend to generate mostly canonical token sequences (Geh et al., 2024), hence, the probabilities $d_{\mathbf{s}}(t)$ generated by an LLM for tokens $t$ that lead to non-canonical sequences ${\mathbf{s}}\shortmid t$ are usually small. More specifically, let $d_{\mathbf{s}}(\text{canonical})$ be the probability mass on the subset of tokens that lead to canonical sequences, i.e., $d_{\mathbf{s}}(\text{canonical})=\sum_{t\in V:\ {\mathbf{s}}\shortmid t\text{ is canonical}}d_{\mathbf{s}}(t)$, then Algorithm 1 requires, in expectation, fewer than $1/d_{\mathbf{s}}(\text{canonical})$ evaluations of canonicity before successfully sampling the next token. That is because, unlike (independent) rejection sampling from $d_{\mathbf{s}}$, which would require in expectation exactly $1/d_{\mathbf{s}}(\text{canonical})$ evaluations of canonicity until a token that leads to a canonical sequence is successfully sampled, our approach never checks the same token twice, which results in an increase in the success probability of sampling a token that leads to a canonical sequence after each failed attempt. [^9] The number of evaluations of canonicity in rejection sampling is distributed according to a geometric distribution with success probability $d_{\mathbf{s}}(\text{canonical})$ resulting in $1/d_{\mathbf{s}}(\text{canonical})$ evaluations in expectation until a successful sample.
The simplest way to test whether the sequence $\mathbf{s}\shortmid t$ is canonical is to compute and check if $\texttt{enc}(\texttt{dec}(\mathbf{s}\shortmid t))=\mathbf{s}\shortmid t$. For BPE-based tokenizers however, it has been shown that it is sufficient to test if $\texttt{enc}(\texttt{dec}(t_{\text{last}}\shortmid t))=t_{\text{last}}\shortmid t$, where $t_{\text{last}}$ is the final token in $\mathbf{s}$. In fact, a recently proposed efficient algorithm to test whether $\mathbf{s}\shortmid t$ is canonical only partially applies the BPE algorithm to $\texttt{dec}(t_{\text{last}}\shortmid t)$ (Vieira et al., 2025b; Hayase et al., 2025).
4.3 Performance of Canonical Generation
Task Metric Llama8B Qwen7B
Standard Canonical Standard Canonical
Translation Quality score $0.72\pm 0.02$ $0.70\pm 0.02$ $0.73\pm 0.01$ $0.71\pm 0.01$
Time per token (s) $0.019$ $0.020$ $0.018$ $0.019$
Non-canonicity rate $6\%$ - $18\%$ -
Spell Checking 11 - edit distance $0.62\pm 0.04$ $0.61\pm 0.04$ $0.74\pm 0.04$ $0.72\pm 0.04$
Time per token (s) $0.020$ $0.023$ $0.018$ $0.018$
Non-canonicity rate $10\%$ - $19\%$ -
Rephrasing Cosine similarity $0.84\pm 0.02$ $0.84\pm 0.02$ $9\pm 0.02 0.88±0$$8\pm 0.02$
Time per token (s) $0.020$ $0.020$ $0.018$ $0.020$
Non-canonicity rate $6\%$ - $5\%$ -
MGSM Accuracy $0.37\pm 0.06$ $0.37\pm 0.06$ $0.63\pm 0.05$ $0.62\pm 0.06$
Time per token (s) $0.020$ $0.021$ $0.018$ $0.020$
Non-canonicity rate $22\%$ - $29\%$ -
Table 1: Performance, (generation) time per token, and non-canonicity rate. The results comprise pairs of outputs generated with standard and canonical generation in German language under the same source of randomness. For the time per token, confidence intervals are not shown, as they were all smaller than $10^{-4}$.
@@ 4.3 Performance of Canonical Generation
We first show that, in comparison with standard generation, the distribution of tokens generated by canonical generation is provably closer to the true distribution of sequences that the LLM has seen during training. Formally, let $p$ denote the true distribution over token sequences ${\mathbf{s}}\in V^{+}$ used during training, for which note that $p(\mathbf{s})=0$ holds for all sequences $\mathbf{s}$ that are non-canonical. Moreover, let $d$ denote the distribution over token sequences that the LLM generates using standard generation, and $\tilde{d}$ the distribution over token sequences that the LLM generates using canonical generation, that is, sampling from the canonicalized next-token distribution $\tilde{d}_{\mathbf{s}}$ given by Eq. 1 at each step of the generation process. Then, the following theorem shows that $p$ is likely to be closer to $\tilde{d}$ than $d$ in terms of KL-divergence, a result independently established in Proposition 3 of Vieira et al. (2025b): [^10] The proof for Theorem 3 can be found in Appendix D.
@@
Let $d$ be absolutely continuous [^11] Absolute continuity is required for the KL-divergence to be well defined, i.e., we require that $d(\mathbf{s})=0$ implies that $p(\mathbf{s})=0$ for all $\mathbf{s}\in V^{+}$. with respect to $p$. Moreover, assume that there exist $\mathbf{s}\in V^{+}$ and $t_{1},t_{2}\in V$ such that $\mathbf{s}\shortmid t_{1}$ is non-canonical with $d(\mathbf{s}\shortmid t_{1})>0$ and $\mathbf{s}\shortmid t_{2}$ is canonical with $p(\mathbf{s}\shortmid t_{2})>0$ and $d(\mathbf{s}\shortmid t_{2})>0$. Then, it holds that
@@
In simpler terms, the two conditions under which canonical generation brings the output token sequences closer to the true distribution are that (i) there exist non-canonical token sequences with positive probability of being generated under $d$ so that their probability mass can be redistributed, and (ii) there exist canonical token sequences with positive probability under $d$ and $p$ so that the redistribution of probability mass in $\tilde{d}$ is beneficial. To understand the intuition behind Theorem 3, note that, by using canonical generation (i.e., sampling from $\tilde{d}$ instead of $d$), the probability that an LLM generates non-canonical token sequences becomes zero, and the probability that it generates any other (canonical) token sequence increases under $\tilde{d}$. Further, since only canonical token sequences have positive probability under the true distribution $p$, this redistribution of probability mass from non-canonical token sequences to canonical ones can only bring the distribution $\tilde{d}$ closer to the true distribution $p$ compared to $d$.
@@
On the flip side, it is important to clarify that a similar property does not necessarily hold for the respective distributions over strings. That is, using canonical generation, the distribution of output strings, resulting from decoding the output token sequences, is not guaranteed to be closer (in terms of KL-divergence) to the true distribution of output strings used during training. Formally, let $p_{\texttt{dec}}=P_{\mathbf{s}\sim p(\mathbf{s})}[\texttt{dec}(\mathbf{s})]$ be the true distribution over strings, $d_{\texttt{dec}}=P_{\mathbf{s}\sim d(\mathbf{s})}[\texttt{dec}(\mathbf{s})]$ be the distribution of strings induced by the distribution of output token sequences $d$, and $\tilde{d}_{\texttt{dec}}=P_{\mathbf{s}\sim\tilde{d}(\mathbf{s})}[\texttt{dec}(\mathbf{s})]$ be the distribution of strings induced by the distribution of output token sequences $\tilde{d}$. Then, we have that
KL $(p_{\texttt{dec}},\tilde{d}_{\texttt{dec}})=$
$=\sum_{\bm{\sigma}\in\Sigma^{+}}p_{\texttt{dec}}(\texttt{dec}(\mathbf{s})=\bm{\sigma})\ln\left(\frac{p_{\texttt{dec}}(\texttt{dec}(\mathbf{s})=\bm{\sigma})}{\tilde{d}_{\texttt{dec}}(\texttt{dec}(\mathbf{s})=\bm{\sigma})}\right)$
$=\sum_{\texttt{enc}(\bm{\sigma}),\bm{\sigma}\in\Sigma^{+}}p(\mathbf{s}=\texttt{enc}(\bm{\sigma}))\ln\left(\frac{p(\mathbf{s}=\texttt{enc}(\bm{\sigma}))}{\tilde{d}(\mathbf{s}=\texttt{enc}(\bm{\sigma}))}\right)$
$=\sum_{\mathbf{s}\in V^{+}:\,\mathbf{s}\text{ is canonical}}p(\mathbf{s}=\mathbf{s})\ln\left(\frac{p(\mathbf{s}=\mathbf{s})}{\tilde{d}(\mathbf{s}=\mathbf{s})}\right)$
[…]
@@
because there is a one-to-one mapping determined by the encoder enc from any string to a canonical token sequence, and only canonical token sequences have positive probability under $p$ and $\tilde{d}$. In contrast, one cannot claim the same for $\textrm{KL}(p_{\texttt{dec}},d_{\texttt{dec}})$ and $\textrm{KL}(p,d)$, as the same string can have multiple tokenizations that have positive probability under $d$. Thus, we cannot conclude that $\textrm{KL}(p_{\texttt{dec}},\tilde{d}_{\texttt{dec}})<\textrm{KL}(p_{\texttt{dec}},d_{\texttt{dec}})$.
Next, given this theoretical gap and since users typically derive value from the string that the output token sequence represents rather than the token sequence itself, we empirically compare the performance and efficiency of canonical and standard generation on the same three tasks from Section 3 (i.e., translation, spell checking, and rephrasing), as well as a standard benchmark for multilingual LLM evaluation (i.e., MGSM), using (open-weights) Llama and Qwen models. [^12] In our experiments, for standard generation, we sample from the next-token distribution $d_{\mathbf{s}}(t)$ using the default categorical sampler in PyTorch, which is an implementation of Gumbel-Max sampling. To this end, we first sample $100$ pairs of outputs per input prompt using standard and canonical generation with the same source of randomness, following Corvelo Benz et al. (2026). Then, we identify the pairs in which the output generated using standard generation are non-canonical, which are the only ones in which standard and canonical generation differ (under Gumbel-Max based sampling), and measure performance and time per token by (re-)sampling $10$ continuations from the token in which the output became non-canonical under standard generation using both standard and canonical generation on each corresponding output (again with the same source of randomness).
To measure performance, we use (i) a quality score provided by a pre-trained neural network for reference-free translation evaluation (Guerreiro et al., 2024) in the translation task, (ii) the (normalized) Levenshtein edit distance (Levenshtein, 1966) between the generated text and the original text without typos in the spell checking task, (iii) cosine similarity of sentence embeddings (Reimers and Gurevych, 2019) of the original and rephrased text in the rephrasing task, and (iv) average accuracy in the MGSM task. Refer to Appendix F for additional details regarding the experimental setup.
Table 1 summarizes the results for the German language, which show that both canonical and standard generation are comparable both in terms of performance and efficiency. Here, the slightly lower performance of canonical generation can be attributed to a limitation shared by constrained generation in general, namely, occasionally restricting the sampling space to low probability generation paths (Vieira et al., 2025b). Refer to Appendix G.2 for qualitatively similar results in other languages.